If the circle with centre(0,m) has equation x^2+y^2+6y+5=0, what is the value of m?
Answer by stanbon(75887)

If the circle with centre(0,m) has equation x^2+y^2+6y+5=0, what is the value of m?
---------------------
x^2+y^2+6y+5=0
Complete the square on the y-terms:
x^2 + y^2+6y+9 = -5+9
(x-0)^2 + (y+3)^2 = 4
This circle has center at (0,-3) and a radius of 2
So m = -3
=======================================================
Cheers,
Stan H.